OpenStudy (anonymous):

write 6y=x+5 in standard form using integers.

OpenStudy (mathmale):

Would you mind writing out the general "equation of a line in standard form" as it's taught in your course? It's either Ax+By=C or Ax+By+C=0. Once you've chosen one of those, we can proceed and answer this question.

OpenStudy (anonymous):

It doesn't say but it's multiple choice if that helps? A) -x+6y=5 B) 6x-6y=30 c) -6x+6y=30 D) -6x-6y=30

OpenStudy (mathmale):

All of these choices have the form Ax+By=C, so that answers my question. Write 6y=x+5 in standard form by simply rearranging the terms so that the equation looks like Ax+By=C.

OpenStudy (anonymous):

I'm not sure how to do that? could you show me please?

OpenStudy (anonymous):

is it -6y+6y=30?

OpenStudy (anonymous):

i meant -6x+6y=30

OpenStudy (mathmale):

This is OK except for the coefficient of x.

OpenStudy (mathmale):

6y=x+5 -x= -x -- ---- -x+6y = 5

OpenStudy (anonymous):

thank you very much for your help!!

OpenStudy (mathmale):

Conetta, my great pleasure!